# Characterizing the MrDeepFakes Sexual Deepfake Marketplace

### v2

@@ 6.4 Hardware
Adequate hardware is chiefly important for creating deepfakes. GPU requirements specified by community members ranged from 4–12 GB of VRAM with recommended GPUs from NVIDIA and AMD. When asked for entry level recommendations, an authoritative member (staff/moderator) recommended an “RTX 3060 minimum or used 1080Ti/2080Ti, nothing else is worth anything for deepfakes (too slow or not enough VRAM), you want 11–12 minimum, that should be plenty until you become really good, then you can get something better with 16-24GB of VRAM.” Hardware costs and electricity costs appear to be prohibitive for some creators:
I wish [I could buy my own GPU] :( Where I live I can’t get anything good and to import the costs will destroy me lol. That’s the dream though one day
@@
Furthermore, cloud GPU users are subject to provider ToS; for instance, in May 2022, Google banned the use of Colab for training deepfakes [google-colab-ban], which immediately caused ripples in the community:
Google banned making deepfakes on Colab for free, if you don’t have Pro or Pro+ plan that’s what happens it will keep disconnecting you, eventually it might even further restrict your google account from being used for Colab.
@@
Honestly? Never used [Runpod], but there should be tutorials on youtube, or try paperspace, but it might be more pricey. If electricity prices keep increasing like that I might soon be forced to use this cloud rubbish too…
@@
Aligned with prior explorations of the efficacy of deplatforming abusive actors [chandrasekharan2017you, levchenko2011trajectories, kanich2008spamalytics, bellingcat2023any] and the pursuit of target hardening [redmiles2024friction], we anecdotally find that while major platform policy changes do not exhaustively eradicate abuse, they do significantly increase the cost of deepfake creation; such changes in the deepfake value chain may be enough to turn away less capable (e.g., financially, technically) creators and stem the growth of this ecosystem.
@@ 6.4.1 Google Colab Deepfake Ban
In May 2022, Google banned using Colab notebooks for training models to generate deepfakes [google-colab-ban]. In the following month, users MrDeepFakes members began to respond with reports of service interruptions, calls for help to the community, and speculations of how the ban was implemented and enforced:
I think they are somehow detecting that the notebook is running DFL processes and terminate it altogether. I really hope this is just temporary. I don’t know of any other free alternatives to Colab for DFL projects.
@@
Members noted that Colab was flagging deepfake-related notebooks by detecting use of the DFL library. Almost immediately, the MrDeepFakes community sought to identify opportunities to circumvent the ban. Several members shared trivial yet effective solutions that obfuscated the source of deepfake code by forking the original DeepFaceLab DFL repository and then cloning the fork, reducing traceability to the banned DeepFaceLab repository:
It doesn’t require any coding knowledge, just copy and paste [the notebook cell changes]. Basically idea is to…circumvent Google ban by changing the cell structure and DFL GitHub source [by forking the repository]. It will take max 30 minutes.
DFL repository:
@@
While discussion of Colab has slowed since the ban, continued reports of service interruptions and iterative development of ban evasion solutions illustrate the cat-and-mouse game between Google and the community. The community concluded that Google’s detection of DeepFaceLab DFL naively relies on a blocklist, resulting in recommendations to “not share [their DFL] Github fork and try to give [it] a unique name.” The risk of informing Google’s deepfake detection has also led to some members seeking to move ban evasion support to private channels, asking others to “please DM if [they] know of any tricks.” As of July 2024, it is unclear whether community solutions still work today; some users report that they are able to use Colab, whereas others report errors.
@@ 6.5 Community Support
To support users in the complex process of creating deepfakes, the MrDeepFakes community provides a place to ask technical questions and receive help. To understand community norms, we analyze $3256$ question threads (i.e., threads placed in the question section) containing $14521$ posts from two “Questions” subsections for both parent sections, “Deepfake Video Creation Tools” and “Stable Diffusion,” to understand the community’s dynamics.
The MrDeepFakes community appears to be responsive and welcoming to newcomers. 80.6% of question threads on MrDeepFakes received a reply, with the median time to receive a reply of 3.4 hours ($\mu=5.8$ days, $\sigma=39.3$ days), which is similar to Stack Overflow, where 80% of posts receive a response and posts have a median response time of 40 minutes ($\mu=11.7$ days, $\sigma=99.2$ days). While MrDeepFakes’s median time to response is longer than that of Stack Overflow, we suspect that this is largely a function of community size; as of August 2024, the Stack Overflow community had 26M members [stackexchange2024sites], as opposed to the 611K members of MrDeepFakes. On the other hand, the mean response time on MrDeepFakes is half that of Stack Overflow, illustrating that the distribution of response times on MrDeepFakes is much tighter. We also find that 72% of questions on MrDeepFakes receive a response in less than three days. Therefore, while members of MrDeepFakes may not receive instantaneous responses, they can reliably expect to receive technical support within a reasonable time frame. Newcomers with accounts under 45 days old start 39.3% of all question threads
Unlike Stack Overflow, which is known to enforce a “no pleases or thank yous” rule [stackexchange2024sites], the MrDeepFakes community does not have any rules regarding the tone or delivery of questions in these forum sections. Anecdotally, we observe that the community encourages pleasantries and more personal connections between members, in part, perhaps because enterprising members are looking to attract clientele:PM me if you’ll need more help, I offer personal courses…Or read my guide, it covers all the basics if you never used DFL to make fakes.
@@
we’re the only non-scam deepfake site on the web and only place with real forum and community, last thing owners want is to scam people
@@
7 Discussionand Conclusion
Grounded in our data, we present the following contributionsDeepfake IBSA causes and amplifies a wide range of real harms. It is commonly used to financially sextort targets, generate “revenge pornography,” and explore sexual fantasies without obtaining consent [wei2024understanding, burkell2019nothing, kshetri2023economics, okolie2023ibsa]. Development of generative AI technology has made deepfake capabilities more accessible, promoting widespread use—including by children [maiberg22clicks]. In this work (Sections 46), we: (1) we measure measured the growth of MrDeepFakes and the targets of deepfake videos; (2) we characterize characterized the participants of the deepfake market and analyze analyzed the prices of bespoke sexual deepfakes; and (3) we document documented the required components of deepfake creation and how the community supports members in the process. We synthesize the implications of our findings on disincentivizing deepfake NCII production and mitigating its harms below.
@@ 7.1 MrDeepFakes as a Community
MrDeepFakes has grown immensely over the past few years, drawing over 1.5 billion views of 43K sexual deepfake videos . as of November 2023. The scale of site engagement highlights that consumption and discussion of sexual deepfakes occurs openly online and has become mainstream . Its and occurs openly online. This growth has led to the maturation of its profit-driven marketplace for sexual deepfakes. As the market matures, we suspect we will observe even more seller specialization than currently exists, continues to develop, seller specialization will likely also further develop similar to other specialized labor markets [sunwoong1989labor]—for . The market will likely also expand. For example, in response to tightening margins on celebrity deepfakes, one MrDeepFakes member suggested that you could make money by applying existing tools to do “many things” beyond celebrity fakes (Section 5.4).
Deepfakes as Art
We also found that the
@@
Deepfakes as Art. The MrDeepFakes community is welcoming and responsive; many of its members and leadership leaders share a value system that positions deepfake creators as artists and the sexual deepfakes they produce as artistic contributions that benefit their the community. This novel incentive unique motivation differentiates deepfake creators as a class of actors separate from previously studied cybercrime and online abuse attackers [thomas2021hatesok, chng2022hacker]; thusactors [thomas2021hatesok, chng2022hacker]. To better understand and disincentivize this class of perpetrator, we identify two immediate three areas for future work: (1) understanding to what degree deepfake creatorsperceive the harms (and the “benefits ”) first, characterizing deepfake creators’ perceptions of the benefits and harms of their workand (2) investigating potential ; second, investigating the applicability of interventions for deepfake creators, as such as targeting inaccurate meta-cognitive (i.e., second-order) beliefs and nudging, which prior work has found effective in for combating misinformation [Jahanbakhsh2024signaling], antisocial behaviors [cheng2017anyone], [jeffrey2021understanding] and poor cybersecurity behaviors [zimmerman2021nudge]. practices [zimmerman2021nudge], respectively; third, drawing on Insoll et al.’s study of child sexual abuse material offenders [insoll2024factors], investigating the prevalence of help-seeking behavior (i.e., wishing to change) among deepfake creators to inform more effective interventions.
Open Source and Academic Work.
Core members are also
@@
Open Source and Academic Work. Core contributors to MrDeepFakes are technically skilled and willing excited to share their technical expertise on deepfake creationexpertise on creating deepfakes. Contrary to the beliefs of some open source developersof a SFW deepfake tool studied by Widder et al.  [widder2022ethicalai], denying technical support does not appear a viable mitigation for the abuse of deepfake creation toolsis not a viable intervention for deepfake tool abuse. Furthermore, significant sharing of academic work within the community shows that its members closely follow and utilize academic research, particularly from AI venues, highlighting the dual-use potential for many computing innovations. Similar to the OSS community, many instances of academic research have used technology neutrality and inevitability to justify the publication and dissemination of innovative developments. We acknowledge that the solution does not lie in silencing research findings or technological innovation, but we argue that researchers and developers should consider principles of “target hardening” in their own work to: (1) carefully choose how to distribute their workand , (2) rigorously evaluate and improve upon existing AI alignment methods, and, when possible, (23) preemptively design and integrate safeguards to promote the intended use and discourage the abuse of their researchtools.
@@ 7.2 Deepfake Abuse Mitigation is Adversarial
Technical solutions to for preventing deepfake creation are will likely be limited, as evidenced by adversarial the successful efforts of MrDeepFakes members to circumvent existing efforts by cloud GPU anti-abuse mechanisms added by cloud providers. While identifying platform-centered choke points (e.g., payment processors, hosting providers, ad providers) can reduce harm, as the security community has concluded for numerous abuse other ecosystems [franklin2007inquiry, han2022infra, krebs2010body], the endless emergence of alternative “bulletproof” service providers and cryptocurrencies has limited the efficacy of policy-centered mitigation (i.e., deplatforming)mitigations like deplatforming. In the face of complex security problems with no tractible, perfect solutions, we acknowledge emphasize the impact and importance of target hardening [redmiles2024friction]; however, we identify also argue that legislation may prove a promising approach to more effectively stem deepfake abuse through legislationfor curbing the creation of non-consensual sexual deepfakes.
Opportunity for Legislation.
Banning Deepfake NCII.A MrDeepFakes staff member shared that the enactment of legislation internationally has already begun to make complicate monetizing sexual deepfakesmore difficult, implying that laws have significant sway over the deepfake economy (Section 5.4). In February 2024, following a push for federal legislation protecting against deepfake abuse [sarnoff2024taylor], forum members asked staff for their reflections on the implications of federally outlawing deepfake abuse, to which a MrDeepFakes staff member replied:
if [abusive deepfakes] get banned then I’ll finally be able to pursue other things in life and won’t feel bad about quitting, also will have more time to work on…other creative work that I can actually show of to people and be proud of making instead of porn
@@
This suggests that laws are powerful enough to disrupt behaviors of Laws may dissuade even highly skilled, core members of the community. We argue that augmenting our existing mitigation approach by incorporating legislative protections has significant potential to more effectively potential to help stem deepfake abuse.
Setting Ethical Norms for Deepfakes.We also observed that MrDeepFakes ’s rules defining celebrities as acceptable targets of deepfake NCII were laxly enforced, indicating a difference between the community’s written rules and normative behaviors in practice. Echoing Timmerman et al’̇s findings [timmerman2023studying], we presented evidence that members’ beliefs regarding the ethics of deepfake NCII creation ranged widely (Section 5.2). Such dissonance even within the MrDeepFakes community illustrates the broader state of flux surrounding the ethics of deepfake content. This liminality presents an opportunity for policy impact beyond raising deepfake creation cost disincentives, through guiding education and legal process approaches. We encourage policymakers to update sexual education curriculum to confront technology-mediated sexual interactions; through early education regarding concepts such as sexual safety, consent, and bodily autonomy in digital-physical ecosystems, we can begin to shape resilient norms surrounding the creation and consumption of deepfake content that proactively mitigate its harms.
Seeking Justice & Reducing Harm.Finally, we preemptively broach legal considerations in light of the mainstreaming of deepfake content and increasing commoditization of deepfake creation tools, as evidenced by the growth of MrDeepFakes and its discussion of limited-data tools. We reiterate Pfefferkorn’s recommendations regarding justice pathways for victims and perpetrators of deepfake NCII: that legislators consider the implications of criminal prosecution for preventing deepfake IBSA—especially regarding children. Due to rising popularity of commodity deepfake NCII tools among teens [maiberg24survey], blanket criminal charges for perpetrators also affect children; Pfefferkorn argues that criminal prosecution of children is counterproductive, unjustifiable, and causes lasting harm [pfefferkorn2024teens]. Therefore, we argue that legislators carefully balance the trade-offs of retributive and restorative justice pathways and acknowledge the nuances of criminalization when children are also known perpetrators.
@@ 8 Conclusion
In this work, we analyzed the MrDeepFakes platform since its inception and characterized the community’s motivations and deepfake creation process. Our investigation surfaces significant growth of deepfake NCII creation and consumption. We further find that beyond nonconsensual sexual gratification and power as understood motivations for IBSA perpetrators, our qualitative analysis reveals novel motivations for buyers of deepfake NCII: the perception of deepfakes as art and the distribution of new deepfakes as a community contribution. We document the requirements for deepfake creation, the challenges it poses, and the respective community-driven solutions. Our work sheds light on the rapidly growing, complex ecosystem around deepfake NCII, and we highlight opportunities to design proactive and reactive interventions to better mitigate deepfake IBSA.
EthicsEthical Considerations
In this paper, we analyze MrDeepFakes, a publicly available platform for hosting, buying, and sellingpublic platform for buying, selling, and hosting sexual deepfakes. Our work is in the spirit of significant prior work by the security community [motoyama11analysis, tseng20tools, wei2024understanding, soska15measuring, gamage2022deepfakes]. However, our context is significantly differentfrom prior workdifferent, and we carefully evaluated the ethical implications of our specific research. We considered many stakeholders when evaluating the ethics of our work: the the potential harm to several stakeholders: the depicted and discussed targets, the MrDeepFakes platformitself, its users, depicted and discussed targets, and the researchers who studied the platform given its potentially harmful contentand the researchers themselves. We discuss the ethical implications of our research choices for each stakeholder below: .
The Platform.

### v3

@@ 6.4 Hardware
Adequate hardware is chiefly important for creating deepfakes. GPU requirements specified by community members ranged from 4–12 GB of VRAM with recommended GPUs from NVIDIA and AMD. When asked for entry level recommendations, an authoritative member (staff/moderator) recommended an “RTX 3060 minimum or used 1080Ti/2080Ti, nothing else is worth anything for deepfakes (too slow or not enough VRAM), you want 11–12 minimum, that should be plenty until you become really good, then you can get something better with 16–24 GB of VRAM.” Hardware costs and electricity costs appear to be prohibitive for some creators:
@@
Furthermore, cloud GPU users are subject to provider ToS; for instance, in May 2022, Google banned the use of Colab for training deepfakes [68], which immediately caused ripples in the community:
@@
Aligned with prior explorations of the efficacy of deplatforming abusive actors [8, 32, 23, 26] and the pursuit of target hardening [55], we anecdotally find that while major platform policy changes do not exhaustively eradicate abuse, they do significantly increase the cost of deepfake creation; such changes in the deepfake value chain may be enough to turn away less capable (e.g., financially, technically) creators and stem the growth of this ecosystem.
@@ 6.4.1 Google Colab Deepfake Ban
In May 2022, Google banned using Colab notebooks for training models to generate deepfakes [68]. In the following month, MrDeepFakes members began to respond with reports of service interruptions, calls for help to the community, and speculations of how the ban was implemented and enforced:
@@
Members noted that Colab was flagging deepfake-related notebooks by detecting use of the DFL library. Almost immediately, the MrDeepFakes community sought to identify opportunities to circumvent the ban. Several members shared trivial yet effective solutions that obfuscated the source of deepfake code by forking the original DFL repository and then cloning the fork, reducing traceability to the banned DFL repository:
@@
While discussion of Colab has slowed since the ban, continued reports of service interruptions and iterative development of ban evasion solutions illustrate the cat-and-mouse game between Google and the community. The community concluded that Google’s detection of DFL naively relies on a blocklist, resulting in recommendations to “not share [their DFL] Github fork and try to give [it] a unique name.” The risk of informing Google’s deepfake detection has also led to some members seeking to move ban evasion support to private channels, asking others to “please DM if [they] know of any tricks.” As of July 2024, it is unclear whether community solutions still work today; some users report that they are able to use Colab, whereas others report errors.
@@ 6.5 Community Support
To support users in the complex process of creating deepfakes, MrDeepFakes provides a place to ask technical questions and receive help from the community. To understand community norms, we analyze $3256$ question threads (i.e., threads in a questions subsection) containing $14521$ posts from two “Questions” subsections for parent sections “Deepfake Video Creation Tools” and “Stable Diffusion.”
The MrDeepFakes community appears responsive and welcoming to newcomers. 80.6% of question threads on MrDeepFakes received a reply, with the median time to receive a reply of 3.4 hours ($\mu=5.8$ days, $\sigma=39.3$ days), which is similar to Stack Overflow, where 80% of posts receive a response and posts have a median response time of 40 minutes ($\mu=11.7$ days, $\sigma=99.2$ days). While MrDeepFakes’ median time to response is longer than that of Stack Overflow, we suspect that this is largely a function of community size; as of August 2024, the Stack Overflow community had 26M members [13], as opposed to the 611K members of MrDeepFakes. On the other hand, the mean response time on MrDeepFakes is half that of Stack Overflow, illustrating that the distribution of response times on MrDeepFakes is much tighter. We also find that 72% of questions on MrDeepFakes receive a response in less than three days. Therefore, while members of MrDeepFakes may not receive instantaneous responses, they can reliably expect to receive technical support within a reasonable time frame. Newcomers with accounts under 45 days old create 39.3% of all question threads
Unlike Stack Overflow, which is known to enforce a “no pleases or thank yous” rule [13], the MrDeepFakes community does not have any rules regarding the tone or delivery of questions in these forum sections. Anecdotally, we observe that the community encourages pleasantries and more personal connections between members, perhaps in part because enterprising members are looking to attract clientele:
@@
7 Discussion
Deepfake IBSA causes and amplifies a wide range of real harms. It is commonly used to financially sextort targets, generate “revenge pornography,” and explore sexual fantasies without obtaining consent [65, 7, 28, 45]. Development of generative AI technology has made deepfake capabilities more accessible, promoting widespread use—including by children [36]. In this work, we: (1) measured the growth of MrDeepFakes and the targets of deepfake videos; (2) characterized the participants of the deepfake market and analyzed the prices of bespoke sexual deepfakes; and (3) documented the required components of deepfake creation and how the community supports members in the process. We synthesize the implications of our findings on disincentivizing deepfake NCII production and mitigating its harms below.
@@ 7.1 MrDeepFakes as a Community
MrDeepFakes has grown immensely over the past few years, drawing over 1.5B views of 43K sexual deepfake videos as of November 2023. The scale of engagement highlights that consumption and discussion of sexual deepfakes has become mainstream and occurs openly online. This growth has led to the maturation of its profit-driven marketplace for sexual deepfakes. As the market continues to develop, we expect seller specialization to evolve in tandem, similar to other specialized labor markets [25]. The market will likely also expand. For example, in response to tightening margins on celebrity deepfakes, one MrDeepFakes member suggested that you could make money by applying existing tools to do “many things” beyond celebrity fakes (Section 5.4).
@@
Deepfakes as Art. The MrDeepFakes community is welcoming and responsive; many of its members and leaders share a value system that positions deepfake creators as artists and the sexual deepfakes they produce as artistic contributions that benefit the community. This unique motivation differentiates deepfake creators from previously studied cybercrime and online abuse actors [61, 9]. To better understand and disincentivize this class of perpetrator, we identify three areas for future work: first, characterizing deepfake creators’ perceptions of the benefits and harms of their work; second, investigating the applicability of interventions for deepfake creators, such as targeting inaccurate meta-cognitive (i.e., second-order) beliefs and nudging, which prior work has found effective for combating misinformation [31] and poor cybersecurity practices [70], respectively; third, drawing on Insoll et al.’s study of child sexual abuse material offenders [21], investigating the prevalence of help-seeking behavior (i.e., wishing to change) among deepfake creators to inform more effective interventions.
@@
Open Source and Academic Work. Core contributors to MrDeepFakes are technically skilled and eager to share their expertise in creating deepfakes. Contrary to the beliefs of some open source developers [67], denying technical support is not a viable intervention for deepfake tool abuse. Furthermore, significant sharing of academic work within the community shows that its members closely follow and utilize academic research, particularly from AI venues, highlighting the dual-use potential for many computing innovations. Similar to the OSS community, many instances of academic research have used technology neutrality and inevitability to justify the publication and dissemination of scientific developments. We acknowledge that the solution does not lie in silencing research findings or technological innovation, but we argue that researchers and developers should consider principles of “target hardening” in their own work to: (1) choose how to distribute their work, (2) rigorously evaluate and improve upon existing AI alignment methods, and, when possible, (3) preemptively design and integrate safeguards to promote the intended use and discourage the abuse of their tools.
@@ 7.2 Deepfake Abuse Mitigation is Adversarial
Technical solutions for preventing deepfake creation will likely be limited, as evidenced by the successful efforts of MrDeepFakes members to circumvent anti-abuse mechanisms added by cloud providers. While identifying platform-centered choke points (e.g., payment processors, hosting providers, ad providers) can reduce harm, as the security community has concluded for numerous other ecosystems [14, 17, 27], the emergence of alternative “bulletproof” service providers and cryptocurrencies has limited the efficacy of policy-centered mitigations like deplatforming. In the face of complex security problems with no tractable, perfect solutions, we emphasize the impact and importance of target hardening [55]; we also argue that legislation may prove a promising approach for curbing the creation of nonconsensual sexual deepfakes.
Banning Deepfake NCII. A MrDeepFakes staff member shared that the enactment of legislation internationally has already begun to complicate monetizing sexual deepfakes, implying that laws have sway over the deepfake economy (Section 5.4). In February 2024, following a push for federal legislation protecting against deepfake abuse [58], forum members asked staff for their reflections on the implications of a federal deepfake ban, to which a MrDeepFakes staff member replied:
@@
Laws may therefore dissuade even highly skilled, core members of the community. We argue that augmenting our existing mitigation approaches by incorporating legislative protections has potential to more effecitvely stem deepfake abuse.
Setting Ethical Norms for Deepfakes. We also observed that MrDeepFakes’ rules defining celebrities as acceptable targets of deepfake NCII were laxly enforced, indicating a difference between the community’s written rules and normative behaviors in practice. Echoing Timmerman et al.’s findings [62], we presented evidence that members’ beliefs regarding the ethics of deepfake NCII creation ranged widely (Section 5.2). Such dissonance even within the MrDeepFakes community illustrates the broader state of flux surrounding the ethics of deepfake content. This liminality presents an opportunity for policy impact beyond raising deepfake creation cost disincentives, through guiding education and legal process approaches. We encourage policymakers to update sexual education curriculum to confront technology-mediated sexual interactions; through early education regarding concepts such as sexual safety, consent, and bodily autonomy in digital-physical ecosystems, we can begin to shape resilient norms surrounding the creation and consumption of deepfake content that proactively mitigate its harms.
Seeking Justice and Reducing Harm. Finally, we preemptively broach legal considerations in light of the mainstreaming of deepfake content and increasing commoditization of deepfake creation tools, as evidenced by the growth of MrDeepFakes and its discussion of limited-data tools. We reiterate Pfefferkorn’s recommendations regarding justice pathways for victims and perpetrators of deepfake NCII: that legislators consider the implications of criminal prosecution for preventing deepfake IBSA—especially regarding children. Due to rising popularity of commodity deepfake NCII tools among teens [37], blanket criminal charges for perpetrators also affect children; Pfefferkorn argues that criminal prosecution of children is counterproductive, unjustifiable, and causes lasting harm [50]. Therefore, we argue that legislators carefully balance the trade-offs of retributive and restorative justice pathways and acknowledge the nuances of criminalization when children are also known perpetrators.
@@ 8 Conclusion
In this work, we analyzed the MrDeepFakes platform since its inception and characterized the community’s motivations and deepfake creation process. Our investigation surfaces significant growth of deepfake NCII creation and consumption. Beyond nonconsensual sexual gratification and power as understood motivations for IBSA perpetrators, our qualitative analysis reveals novel motivations for buyers of deepfake NCII: the perception of deepfakes as art and the distribution of new custom deepfakes as a community contribution. We document the requirements for deepfake creation, the challenges it poses, and the respective community-driven solutions. Our work sheds light on the rapidly growing, complex ecosystem of deepfake NCII, and we highlight opportunities to design proactive and reactive interventions to better mitigate deepfake IBSA.
Acknowledgments
The authors thank Riana Pfefferkorn, Hans Hanley, and the anonymous reviewers for their feedback and suggestions. This work was supported in part by a Sloan Research Fellowship, a Magic Grant from the Brown Institute for Media Innovation, an NSF Graduate Research Fellowship #DGE-1656518, and by the National Science Foundation under Grant Numbers #2030859 and #2127309 to the Computing Research Association for the CIFellows Project. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
Ethical Considerations
In this paper, we analyze MrDeepFakes, a public platform for buying, selling, and hosting sexual deepfakes. Our work is in the spirit of prior work by the security community [41, 63, 65, 60, 15]. However, our context is different, and we carefully evaluated the ethical implications of our specific research. We considered the potential harm to several stakeholders: the depicted and discussed targets, the MrDeepFakes platform, its users, and the researchers themselves. We discuss the ethical implications for each stakeholder below.